OpenStudy (anonymous):

hey can anybody solve this question .... if f(x)=x2-x,show that f(x+1)=f(-x)

OpenStudy (anonymous):

f(x+1) = (x+1)^2 - (x+1) = x^2 +2x +1 - x -1 =x^2 +x = f(-x)

OpenStudy (anonymous):

f(x+1)=(x+1)^2 -(x+1) = (x+1)((x+1-1) = x(x-1)=f(x)...ur statement is wrong..maybe??

OpenStudy (anonymous):

hey thanx mate

OpenStudy (anonymous):

but iam still blank :S

OpenStudy (anonymous):

nah ive checked it statement is right

OpenStudy (anonymous):

ankur \[f(x)\neq x(x-1)\]

OpenStudy (anonymous):

err nvm, durr

OpenStudy (radar):

\[f(x)=x ^{2}-x\] This was given Show that f(x+1) = f(-x). Work out each one and see if it is true First do f(x+1)\[f(x+1)=(x+1)^{2}-(x+1)=x ^{2}+2x+1-x-1=x ^{2}+x\] Now work it out for -x \[f(-x)=(-x)^{2}-(-x)=x ^{2}+x\] The results look the same to me